OpenStudy (anonymous):

What is the slope of the line whose equation is 12 = 4x – 6y?

OpenStudy (anonymous):

To find the y-intercept, you have to let X = 0... Therefore you get 4(0)-6y=12 -6y=12 y = -2 To get the slope, you have to rearrange the equation into y = mx + c form. Therefore. -6y = -4x + 12 You need a single y, so divide through by -6. y = -4x/-6 +12/-6 Your slope is the number in front of the x variable. I.e -4/-6 Simplify it to, 2/3...